I've tried really hard with this question but I just don't get it, I don't really get any algebra.
Jamie and Phillip were testing two different cars and in a given amount of time found that Jamie had travelled three times as far as Phillip. If the total distance is 154km, how many more kilometres did Jamie travel than Phillip?

Answer by mananth(16949)

Let Philip travel x km
Jamie travels 3x
..
3x+x=4x=154
4x=154
/4
x=38.5 km ---Philip
3x= 3*38.5= 115.5--- Jamie
Find the difference .